Desribe or calculate the following things of Microsoft corporation from 2019 (data is presented in the Microsoft corporation 10-k report 2019)

Operating expenses

Tax expense

Total Dividends

Cash to/from operating activities

Cash to/from investing activities

Cash to/from financing activities

Net change in cash
Answer #1

1. Operating expenses = Research and development expenses + Sales and marketing expenses + General and administrative expenses

= 16,876 +18213+ 4885

= 39974 million

2. Current tax expense = U.S. federal + U.S. state and local + Foreign = 4718+ 662+5531 = 10911 million

3. Total Dividends = 14107 million

4. Cash to/from operating activities = 52,185 million

5. Cash to/from investing activities = ( 15773) million

6. Cash to/from financing activities = (36887) million

7. Net change in cash = (590) million

10K: microsoft.com/investor/reports/ar19/download-center/